OpenStudy (anonymous):

What is the solution to this equation? log2(3x-1)=log4(x+8)

OpenStudy (anonymous):

change of base would be useful

OpenStudy (anonymous):

to solve?

OpenStudy (anonymous):

yes \[\large log_2(a)=log_4(b)\] \[\frac{ln(a)}{ln(2)}=\frac{ln(b)}{ln(4)}\]

OpenStudy (anonymous):

\[\frac{ln(a)}{ln(2)}=\frac{ln(b)}{ln(4)}\] \[ln(a)=\frac{ln(2)}{ln(4)}ln(b)\] \[ln(a)=k~ln(b)\] \[ln(a)=ln(b^k)\] \[a=b^k\]

OpenStudy (anonymous):

ln4/ln2 might be more practical: since log2(4) = 2, otherwise that k is a sqrt

OpenStudy (anonymous):

it simply reduces to some quadratic form if followed (3x-1)^2 = x+8